OpenStudy (anonymous):

find the zeros of the function f(x) = x^2+6x+18

OpenStudy (anonymous):

\[f(x)=x^2+6x+18\] \[x^2+6x+18=0\] \[\Delta=b^2-4*a*c\] \[\Delta=6^2-4*1*18\] \[\Delta=-36\] than we have to change the domain \[\mathbb{R}\to\mathbb{C}\] \[i^2=-1\] therefore\[-36=36i^2\] \[x=\frac{-b\pm\sqrt{\Delta}}{2a}\] \[x=\frac{-6\pm\sqrt{36i^2}}{2}\] \[x=\frac{-6\pm6i}{2}\] \[x=-3\pm3i\] \[\boxed{\boxed{x_1=-3+3i~~and~~x_2=-3-3i}}\]

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

nop